OpenStudy (anonymous):

Write the coordinate point for the vertex of this parabola: x=-1/8y^2.

OpenStudy (ranga):

This is a sideways parabola. The vertex form of a sideways parabola is: x = a(y - k)^2 + h where (h,k) is the vertex. Compare x = -1/8y^2 to the vertex form and you can see a = -1/8, h = 0, k = 0 Therefore, the vertex is (0,0) (the origin) a is negative and therefore the parabola opens up in the negative x direction.

OpenStudy (anonymous):

thank you :)